Given three floating-point numbers x, y, and z, output x to the power of z, x to the power of (y to the power of z), the absolute value of (x minus y), and the square root of (x to the power of z).

Output each floating-point value with two digits after the decimal point, which can be achieved as follows:
print('{:.2f} {:.2f} {:.2f} {:.2f}'.format(your_value1, your_value2, your_value3, your_value4))

Ex: If the input is:

5.0
1.5
3.2

Then the output is:

172.47 361.66 3.50 13.13
Answer #1

import math

x = float(input("Enter 1st number: "))

y = float(input("Enter 2nd number: "))

z = float(input("Enter 3rd number: "))

x_to_the_power_of_z = pow(x, z);

x_to_the_power_of_y_to_the_power_of_z = pow(x, pow(y, z))

absolute_value_of_x_minus_y = abs(x - y)

square_root_of_x_to_the_power_of_z = math.sqrt(x_to_the_power_of_z)

print('{:.2f} {:.2f} {:.2f} {:.2f}'.format(x_to_the_power_of_z, x_to_the_power_of_y_to_the_power_of_z, absolute_value_of_x_minus_y, square_root_of_x_to_the_power_of_z))

******************************************************************** SCREENSHOT *******************************************************